RAITH at Stanford OK

James Conway toecutter4ranger at gmail.com
Tue Sep 22 14:42:44 PDT 2009


Hi Cole,
As always, it was my pleasure to speak with you this morning.

Status of Stanford RAITH -- ALL OK emission returning to normal levels.

There was an error at 20:22 in the leolog.log file of an extractor set point
trip event. ( Error 519 and not the 818 error code that was originally
reported to me last night. Likely the number was transposed in translation
as the User was upset by this event.)

Upon coming in this morning, I observed excellent FE-Gun and SYSTEM vacuum.
(5.7E-010 mbar GUN and 7.7e-007 SYSTEM)  However the Extraction I for the
filament was very low ~ 44 uA .
I checked the gun parameters and identified that the Extractor Voltage has
not come all the way up after the previous trip to its previous setponi
t.  Extraction Voltage when ramping up initially had stopped at 4.77 kV
instead of the voltage reaching the default setting now at 5.60 kV.

Emission Current will be near optimal in a few days.  Stay tuned for further
updates.  Beam I is stable and all Users should come back to this system for
the r EBL needs.


Best,

James Conway
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/raith-pcs/attachments/20090922/d23209af/attachment.html>


More information about the raith-pcs mailing list